I need to make a circle by only knowing the circumference has to be only of inches, for example I need a circle that has a circumference of exactly 200 inches, how might I figure this out?
1
Expert's answer
2013-07-03T09:02:38-0400
Thecircumference of a circle is the edge or rim of a circle itself. It is the
equivalent of 'perimeter' for a circle. In other words, if you took a circle
and unrolled it, you would have its circumference. You can use either of the
formulas below to find the circumference . One formula use the radius of a circle;
the other formula uses the diameter.

Formula using Diameter
Circumference=π⋅diamete
Formula using Radius
Circumference=2π⋅radius

These two formulas are just two different ways of finding the same thing
(circumference) because the diameter of a circle =2• radius. So you can use
whichever formula is more convenient. If you know a circle's radius, use the
first version of the formula; if you know the diameter, use the second.
Example:
Circumference=2π⋅radius
200 = 2 * 3.14 * radius
radius = 200/6.28 = 31.84 inches

The Radius is the distance from the center to the edge.
So now, a circle is easy to make:
Draw a curve that is "radius" away from a central point.
And so:
All points are the same distance from the center.